OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair. x = 2y - 2 and 3x - y = 14 (6, 4) (4, 6) (1, 0) (0, 1) no solution

OpenStudy (anonymous):

ans is (6,4)? correct?

OpenStudy (abb0t):

Substitute your x = 2y-2 for x in the second equation. and solve for y.

OpenStudy (abb0t):

Once you have your y, plug in that value back for the first equation to get your x.

OpenStudy (anonymous):

3(2y-2)-y=14?

OpenStudy (anonymous):

3(2y - 2) - y = 14 5y-6 = 14 y=4 => x=6 (6,4)

OpenStudy (anonymous):

how did you get 5y? it would be 6y

OpenStudy (anonymous):

yes but them you have a -y, so 5y

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

6y -6 -y =14

OpenStudy (anonymous):

5y-6=14 i get that but what did you do after that?

OpenStudy (anonymous):

5y = 14+6 5y = 20 y=4 them from 3x - y = 14 3x - (4) = 14 3x = 18 x=6 so (6,4)

OpenStudy (anonymous):

oh ok thank you i get it:)

OpenStudy (anonymous):

any time xD